# Discrete Linear Canonical Transform on Graphs: Uncertainty Principle and Sampling

### v2

@@ 2.3 Sampling on Graph Signals
Consider the sampling set $\mathcal{S}=\left(\mathcal{S}_{0},\dots,\mathcal{S}_{|\mathcal{S}|-1}\right),\mathcal{S}_{i}\in\{0,1,\dots,N-1\}$, its dimension coefficient is defined as $|\mathcal{S}|$. This coefficient is the size of the sampling indexes for obtaining a sampled signal $\mathbf{x}_{\mathcal{S}}\in\mathbb{C}^{|\mathcal{S}|}(|\mathcal{S}|<N)$ from a bandlimited graph signal $\mathbf{x}\in\mathbb{C}^{N}$. The sampling operator $\mathbf{D}$ is defined as a linear mapping from $\mathbb{C}^{N}$ to $\mathbb{C}^{|\mathcal{S}|}$ [11] , which is expressed as
@@
We then recover $\mathbf{x}$ from $\mathbf{x}_{\mathcal{S}}$ with interpolation operator $\mathbf{R}$, which is a linear mapping from $\mathbb{C}^{|\mathcal{S}|}$ to $\mathbb{C}^{N}$.
$\begin{array}[]{rl}\mathrm{sampling:}&\mathbf{x}_{\mathcal{S}}=\mathbf{D}\mathbf{x}\in\mathbb{C}^{|\mathcal{S}|},\\ \mathrm{interpolation:}&\mathbf{x}_{\mathcal{R}}=\mathbf{Rx}_{\mathcal{S}}=\mathbf{RDx}\in\mathbb{C}^{N},\end{array}$
where $\mathbf{x}_{\mathcal{R}}\in\mathbb{C}^{N}$ recovers $\mathbf{x}$ either approximately or exactly.
@@ 3 Uncertainty Principle of the GLCT
In the realm of continuous-time signals, Landau, Pollak, and Slepian extensively explored space-frequency analysis linked to projection operators. Their groundbreaking work, spanning the 1960s, encompassed the uncertainty principle, eigenvalue distributions, and prolate ellipsoid wave functions [32, 33]. By extending this uncertainty principle to graph signals [12], we circumvent the need for any extra distance definition, mitigating the drawbacks associated with graph distance definitions. Furthermore, we apply this methodology to establish the uncertainty principle in the GLCT domain.
@@ 3.1 Localization in Vertex and Spectral Domains
3.2 Uncertainty Principle of the GLCT
Using the energy expectations of equations (9) and (10), we can extend them to the GLCT domain similar to [12, 22].
Suppose a vector $\mathbf{x}$, vertex subset $\mathcal{S}$ and graph spectral subset $\mathcal{F}$, two subsets are used (12) and (13). The vectors $\mathbf{Dx}$ and $\mathbf{B}^{\mathbf{M}}\mathbf{x}$ represent the projection of $\mathbf{x}$ on the vertex set $\mathcal{S}$ and the spectral set $\mathcal{F}$ respectively. Then, similar to equations (9) and (10), $\zeta^{2}$ and $(\eta^{\mathbf{M}})^{2}$ are used to denote the energy expectations falling within the sets $\mathcal{S}$ and $\mathcal{F}$ respectively, as follows
@@ 3.2 Uncertainty Principle of the GLCT
Fig. 2 illustrates the uncertainty principle of Theorem 1, highlighting some noteworthy characteristics. The blue portion in the figure corresponds to the range of $\Gamma$ in the GFT domain, while in the GLCT domain, the ranges in the lower left and upper right respectively tend to $(0,0)$ and $(1,1)$ as the parameter continuously changes, as indicated by the additional red portion. If we set the inequalities given by Theorem 1 to equality, we can determine the equations of the curve at the four corners of Fig. 2, namely the upper right, upper left, lower right, and lower left corners. Notably, the upper right corner of the admissible region $\Gamma$ specifies the domain where the maximum energy concentration occurs in the vertex and graph spectral for $\phi(\zeta,\eta^{\mathbf{M}})$. The equation for this curve is obtained
@@ 4 Sampling Theory of the GLCT
In this section, we present a novel sampling theory for the GLCT, drawing inspiration from the established sampling theories of the GFT and the GFRFT [11, 12, 13, 14, 15, 34]. Our approach to graph sampling, rooted in the graph linear canonical basis and graph frequency-based sampling methods, tackles the challenge of reconstructing bandlimited graph signals from their subsampled counterparts.
An $\mathbf{M}$-bandlimited graph signal is characterized as a signal with zero graph linear canonical coefficients aligned with the eigenvalues of $\lambda_{|\mathcal{F}|-1}$. This definition implies that $f(i)=0$ for $\lambda,i\geq|\mathcal{F}|$, where $\mathcal{F}$ denotes the set of indices associated with non-zero graph linear canonical coefficients.
Specifically, the indices in $\mathcal{F}$ identify those components of the graph linear canonical coefficients that contribute significantly to the signal. We further elaborate on the concept of $\mathbf{M}$-bandlimited graph signals in the subsequent discussion.
@@ 4.2 Sampling and Perfect Recovery
Hence, the matrix $\mathbf{R=\left(I-\bar{D}B^{M}\right)}^{-1}$. The perfect recoverability of $\mathbf{x}$ implies the existence of $\mathbf{R}$, which is tantamount to the invertibility of $\mathbf{\left(I-\bar{D}B^{M}\right)}$. Consequently, we infer that when $\left|\left|\mathbf{\bar{D}B^{M}}\right|\right|<1$, the original signal can be perfectly recovered, as reported in [12]. Herein, the maximum norm of $\mathbf{\bar{D}B^{M}}$ being $1$ restricts our consideration to this scenario, where perfect localization on the complement of the sampling set prevails, rendering signal recovery unattainable. This is closely related to the uncertainty principles discussed in Section 3. Furthermore, since $\mathbf{x}\in\mathrm{BL}_{|\mathcal{F}|}\left(\mathbf{O}^{\mathbf{M}}\right)$,
@@
Let $\mathbf{J=RD}$, we have
thus $\mathbf{J=PD}$ is a projection operator.
Since $\mathbf{RD}$ is a projection operator, $\mathbf{RDx}$ is an approximation of $\mathbf{x}$ in the space of $\mathrm{BL}_{|\mathcal{F}|}\left(\mathbf{O}^{\mathbf{M}}\right)$. We achieve perfect recovery when $\mathbf{x}$ is in the space of $\mathrm{BL}_{|\mathcal{F}|}\left(\mathbf{O}^{\mathbf{M}}\right)$. ∎
@@ 4.3 Qualified Sampling Operator
where $\mathbf{A^{M}}$ lacks physical significance as it represents a complex matrix, specifically the adjacency matrix. However, when $\mathbf{M}=(0,1;-1,0)$, $\mathbf{O}^{\mathbf{M}}$ degenerates into $\mathbf{V}^{-1}$, and in this case, $\mathbf{A^{M}}$ serves as the shift operator.
@@
As observed, the sampled signal $\mathbf{x}_{\mathcal{S}}$ and the frequency signal $\mathbf{\hat{x}}_{|\mathcal{F}|}$ in the dual domain form a graph linear canonical operator. Therefore, we obtain a new adjacency matrix, leading to the following theorem similar to [11].
@@ 4.4 Optimal Sampling Operator
When sampling graph signals, aside from determining the appropriate sample size, devising a strategy for selecting sampling node locations is crucial as the positions of the sampled nodes play a pivotal role in the performance of the reconstruction algorithm. Given the multiple choices of $|\mathcal{F}|$ linearly independent rows in $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$, our objective is to opt for an optimal set that minimizes the impact of noise. One strategy involves selecting sampling positions to minimize the normalized MSE (NMSE) [12, 34]. Consider the noise $e$ introduced during the sampling process
@@
In the following, we provide several alternative strategies for selecting sampling sets in the GLCT domain based on the optimal design approaches proposed by [11, 12, 13, 14, 38].
  1. 1.
    Minimization of the Frobenius Norm of $\mathbf{P}$ (MinFro)
    In this strategy, the objective is to directly minimize the Frobenius norm of the matrix $\mathbf{P}$, aiming to minimize the NMSE between the original and recovery signals. The method involves selecting columns from the matrix $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$ to minimize the Frobenius norm of the pseudo-inverse matrix $\mathbf{P}$,
@@
  2. 2.
    Maximization of the Volume of the Parallelepiped Formed With the Columns of $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$ (MaxVol)
    This strategy, inspired by [12], aims to select a set $\mathcal{S}$ of columns from the matrix $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$ to maximize the volume of the parallelepiped constructed with the chosen columns. Essentially, this approach involves completing the matrix $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$. The volume can be computed as the determinant of the matrix $\mathbf{O}^{\mathbf{M}}_{|\mathcal{F}|}\mathbf{D}\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$.
@@
  3. 3.
    Minimization of the singular value of $\mathbf{P}$ (MinPinv)
    Due to $\lambda_{i}\left(\mathbf{B^{M}DB^{M}}\right)=\sigma^{2}_{i}\left(\mathbf{B^{M}D}\right)=\sigma^{2}_{i}\left(\mathbf{\Sigma_{\mathcal{F}}O^{M}D}\right)=\sigma^{2}_{i}\left(\mathbf{P}\right)$, in the case of uncorrelated noise, this is equivalent to minimizing $\sum^{\left|F\right|}_{i}\sigma^{2}_{i}$,
@@
  4. 4.
    Maximization of the minimum singular value of $\mathbf{D}\mathbf{O}^{-\mathrm{M}}_{|\mathcal{F}|}$ (MaxSigMin)
    The fourth strategy is designed to leverage the 2-norm to minimize NMSE, which is equivalent to minimizing the largest singular value of $\mathbf{P}$. In other words, we aim to maximize the smallest singular value $\sigma_{\mathrm{min}}$ of $\mathbf{D}\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$ for every qualified $\mathbf{D}$,
@@
  5. 5.
    Maximization of the singular value of $\mathbf{\mathbf{D}\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}}$ (MaxSig)
    The third strategy aims to select columns of the matrix $\mathbf{O}^{-\mathrm{M}}_{|\mathcal{F}|}$ to maximize its singular values. Although this strategy is not directly related to NMSE optimization, it is easy to implement and demonstrates good performance in practice,
@@
To address the sampling strategies outlined above, we introduce five greedy algorithms aimed at resolving this selection challenge to ascertain the optimal sampling operator. In tackling similar matrix approximation optimization issues, greedy algorithms have demonstrated efficacy in closely approximating the global optimum [39]. They provide an efficient and relatively straightforward approach to resolving the sampling set selection dilemma. By iteratively selecting locally optimal solutions, greedy algorithms approximate the global optimum, thereby achieving a balance between computational complexity and the quality of the solution. Since strategies 1) and 5), as well as 3) and 4), are essentially equivalent, we present Algorithm 1, which summarizes the three sampling strategies 2), 4), and 5). Furthermore, Table 1 encapsulates the A, D, E, T-optimal designs [38] and objective functions associated with the aforementioned sampling strategies.
  Input
           $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$ the first $|\mathcal{F}|$ columns of $\mathbf{O}^{-\mathbf{M}}$
           $|\mathcal{S}|$ the number of samples
           $N$ the number of rows in $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$
  Output
           $\mathcal{S}$ sampling set
  Function
           $\begin{array}[]{l}\mathbf{for~{}}k=1:|\mathcal{S}|\\ \quad s=\arg\max_{j\in\{1:N\}-\mathcal{S}}\sum^{\left|\mathcal{F}\right|}_{i=1}\sigma^{2}_{i}\left(\left(\mathbf{O^{-M}_{|\mathcal{F}|}}\right)_{\mathcal{S}\cup\left\{j\right\}}\right)\\ \text{or}\ s=\arg\max_{j\in\{1:N\}-\mathcal{S}}\sigma_{\mathrm{min}}\left(\left(\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}\right)_{\mathcal{S}+\{j\}}\right)\\ \text{or}\ s=\arg\max_{j\in\{1:N\}-\mathcal{S}}\prod^{|\mathcal{F}|}_{i=1}\lambda_{i}\left(\left(\left(\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}\right)^{\ast}_{S\cup\left\{j\right\}}\left(\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}\right)_{S\cup\left\{j\right\}}\right)\right)\\ \quad\mathcal{S}=\mathcal{S}+\{s\}\\ \mathbf{end}\\ \mathbf{return}~{}\mathcal{S}\end{array}$
@@
We compared our proposed five sampling strategies with random sampling under the GLCT sampling framework. As shown in Fig. 3, the NMSE is reported, defined as the error between the original and recovered signals divided by the original signal for each node. (a) is based on a cycle graph with N = 32 nodes, and the signal is a $\mathbf{M}$-bandlimited signal with parameters $\alpha=0.8,\beta=32,\xi=0.5k+1,k=1,...,N$. The frequency bandwidth is $|\mathcal{F}|=4$. (b) is based on a swiss roll graph with N = 256 nodes, and the signal is a $\mathbf{M}$-bandlimited signal with same parameters. The frequency bandwidth is $|\mathcal{F}|=32$. We compare the NMSE of six sampling methods as the number of sampled nodes increases for two different graph signals. We observed that, with a small number of sampling nodes, all methods except random sampling could effectively recover the signal.

### v3

@@ 2.3 Sampling on Graph Signals
Consider the sampling set $\mathcal{S}=\left(\mathcal{S}_{0},\dots,\mathcal{S}_{|\mathcal{S}|-1}\right),\mathcal{S}_{i}\in\{0,1,\dots,N-1\}$, its dimension coefficient is defined as $|\mathcal{S}|$. This coefficient is the size of the sampling indexes for obtaining a sampled signal $\mathbf{x}_{\mathcal{S}}\in\mathbb{C}^{|\mathcal{S}|}(|\mathcal{S}|<N)$ from a bandlimited graph signal $\mathbf{x}\in\mathbb{C}^{N}$. The sampling operator $\mathbf{D}$ is defined as a linear mapping from $\mathbb{C}^{N}$ to $\mathbb{C}^{|\mathcal{S}|}$, which is expressed as
@@
We then recover $\mathbf{x}$ from $\mathbf{x}_{\mathcal{S}}$ with interpolation operator $\mathbf{R}$, which is a linear mapping from $\mathbb{C}^{|\mathcal{S}|}$ to $\mathbb{C}^{N}$. Sampling is denoted as $\mathbf{x}_{\mathcal{S}}=\mathbf{D}\mathbf{x}\in\mathbb{C}^{|\mathcal{S}|}$, and interpolation is represented by $\mathbf{x}_{\mathcal{R}}=\mathbf{Rx}_{\mathcal{S}}=\mathbf{RDx}\in\mathbb{C}^{N}$, where $\mathbf{x}_{\mathcal{R}}$ recovers $\mathbf{x}$ either approximately or exactly.
@@ 3 Uncertainty Principle of the GLCT
In the realm of continuous-time signals, Landau, Pollak, and Slepian extensively explored space-frequency analysis linked to projection operators. Their groundbreaking work, spanning the 1960s, encompassed the uncertainty principle, eigenvalue distributions, and prolate ellipsoid wave functions [40, 41]. By extending this uncertainty principle to graph signals, we circumvent the need for any extra distance definition, mitigating the drawbacks associated with graph distance definitions. Furthermore, we apply this methodology to establish the uncertainty principle in the GLCT domain.
@@ 3.1 Localization in Vertex and Spectral Domains
3.2 Uncertainty Principle in the GLCT domain
Using the energy expectations of Eq. (8), we can extend them to the GLCT domain similar to [12, 11].
Suppose a vector $\mathbf{x}$, vertex subset $\mathcal{S}$ and graph spectral subset $\mathcal{F}$, two subsets are used (10) and (11). The vectors $\mathbf{Dx}$ and $\mathbf{B}^{\mathbf{M}}\mathbf{x}$ represent the projection of $\mathbf{x}$ on the vertex set $\mathcal{S}$ and the spectral set $\mathcal{F}$ respectively. Then, similar to Eq. (8), $\zeta^{2}$ and $(\eta^{\mathbf{M}})^{2}$ are used to denote the energy expectations falling within the sets $\mathcal{S}$ and $\mathcal{F}$ respectively, as follows
@@ 3.2 Uncertainty Principle in the GLCT domain
Fig. 4 illustrates the uncertainty principle of Theorem 1, highlighting some noteworthy characteristics. The blue portion in the figure corresponds to the range of $\Gamma$ in the GFT domain, while in the GLCT domain, the ranges in the lower left and upper right respectively tend to $(0,0)$ and $(1,1)$ as the parameter continuously changes, as indicated by the additional red portion. If we alter the parameters of the GLCT, its range can either expand or shrink, which contrasts sharply with the fixed range of the GFT. And if we set the inequalities given by Theorem 1 to equality, we can determine the equations of the curve at the four corners of Fig. 4, namely the upper right, upper left, lower right, and lower left corners. Notably, the upper right corner of the admissible region $\Gamma$ specifies the domain where the maximum energy concentration occurs in the vertex and graph spectral for $\phi(\zeta,\eta^{\mathbf{M}})$. The equation for this curve is obtained
@@ 4 Sampling Theory of the GLCT
In this section, we present a novel sampling theory for the GLCT, drawing inspiration from the established sampling theories of the GFT and the GFRFT [12, 13, 14, 15, 16, 17, 18, 19]. Our approach to graph sampling, rooted in the graph linear canonical basis and graph frequency-based sampling methods, tackles the challenge of reconstructing bandlimited graph signals from their subsampled counterparts.
An $\mathbf{M}$-bandlimited graph signal is characterized as a signal with zero graph linear canonical coefficients aligned with the eigenvalues of $\lambda_{|\mathcal{F}|-1}$. This definition implies that $f(i)=0$ for $\lambda,i\geq|\mathcal{F}|$, where $\mathcal{F}$ denotes the set of indices associated with non-zero graph linear canonical coefficients. Specifically, the indices in $\mathcal{F}$ identify those components of the graph linear canonical coefficients that contribute significantly to the signal. We further elaborate on the concept of $\mathbf{M}$-bandlimited graph signals in the subsequent discussion.
@@ 4.2 Sampling and Perfect Recovery
Hence, the matrix $\mathbf{R=\left(I-\bar{D}B^{M}\right)}^{-1}$. The perfect recoverability of $\mathbf{x}$ implies the existence of $\mathbf{R}$, which is tantamount to the invertibility of $\mathbf{\left(I-\bar{D}B^{M}\right)}$. Consequently, we infer that when $\left|\left|\mathbf{\bar{D}B^{M}}\right|\right|<1$, the original signal can be perfectly recovered. Herein, the maximum norm of $\mathbf{\bar{D}B^{M}}$ being $1$ restricts our consideration to this scenario, where perfect localization on the complement of the sampling set prevails, rendering signal recovery unattainable. This is closely related to the uncertainty principles discussed in Section 3. Furthermore, since $\mathbf{x}\in\mathrm{BL}_{|\mathcal{F}|}\left(\mathbf{O}^{\mathbf{M}}\right)$,
@@
The proof is reported in C. ∎
@@ 4.3 Qualified Sampling Operator
where $\mathbf{A^{M}}$ lacks physical significance as it represents a complex matrix, specifically the adjacency matrix. However, when $\mathbf{M}=[0,1;-1,0]$, $\mathbf{O}^{\mathbf{M}}$ degenerates into $\mathbf{V}^{-1}$, and in this case, $\mathbf{A^{M}}$ serves as the shift operator.
@@
As observed, the sampled signal $\mathbf{x}_{\mathcal{S}}$ and the frequency signal $\mathbf{\hat{x}}_{|\mathcal{F}|}$ in the dual domain form a GLCT operator. Therefore, we obtain a new adjacency matrix, leading to the following theorem similar to [13].
@@ 4.4 Optimal Sampling Operator
When sampling graph signals, aside from determining the appropriate sample size, devising a strategy for selecting sampling node locations is crucial as the positions of the sampled nodes play a pivotal role in the performance of the reconstruction algorithm. Given the multiple choices of $|\mathcal{F}|$ linearly independent rows in $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$, our objective is to opt for an optimal set that minimizes the impact of noise. One strategy involves selecting sampling positions to minimize the normalized MSE (NMSE) [12, 13, 14, 15, 16, 17, 18, 19]. Consider the noise $e$ introduced during the sampling process
@@
In the following, we provide several alternative strategies for selecting sampling sets in the GLCT domain based on the A, D, E, T-optimal design approaches proposed by [12, 13, 14, 15, 44].
1) Minimization of the Frobenius Norm of $\mathbf{P}$ (MinFro): In this strategy, the objective is to directly minimize the Frobenius norm of the matrix $\mathbf{P}$, aiming to minimize the NMSE between the original and recovery signals. The method involves selecting columns from the matrix $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$ to minimize the Frobenius norm of the pseudo-inverse matrix $\mathbf{P}$, which can reduce the impact of noise in the restored signal, thereby improving the energy concentration of the signal,
@@
2) Maximizing signal energy concentration in the vertex domain (MaxVertex): This approach aims to select sampling nodes that capture the maximum signal energy, ensuring that the signal energy is maximized at the sampling nodes,
$\mathbf{D}^{\mathrm{opt}}=\arg\max_{\mathbf{D}}\sum^{\mathcal{S}}_{i=1}\left|\mathbf{x}_{i}\right|^{2}=\arg\max_{\mathbf{D}}~{}\left|\mathbf{Dx}\right|^{2}.$ (27)
3) Maximizing signal energy concentration in the spectral domain (MaxSpec): This method selects nodes in the spectral domain that capture the maximum signal energy, i.e., the set of nodes with the highest spectral energy, ensuring optimal concentration of signal at sampling points,
$\mathbf{D}^{\mathrm{opt}}=\arg\max_{\mathbf{D}}~{}\sum^{\left|\mathcal{S}\right|}_{i=1}\left|\left(\mathbf{O^{M}x}\right)_{i}\right|^{2}=\arg\max_{\mathbf{D}}~{}\left|\mathbf{O^{M}Dx}\right|^{2}.$ (28)
4) Maximization of the Volume of the Parallelepiped Formed With the Columns of $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$ (MaxVol): This strategy aims to select a set $\mathcal{S}$ of columns from the matrix $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$ to maximize the volume of the parallelepiped constructed with the chosen columns. Essentially, this approach involves completing the matrix $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$. The volume can be computed as the determinant of the matrix $\mathbf{O}^{\mathbf{M}}_{|\mathcal{F}|}\mathbf{D}\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$, which will improve the condition number of the sampling matrix and thus enhance the energy concentration of the signal,
@@
5) Minimization of the singular value of $\mathbf{P}$ (MinPinv): Due to $\lambda_{i}\left(\mathbf{B^{M}DB^{M}}\right)=\sigma^{2}_{i}\left(\mathbf{B^{M}D}\right)=\sigma^{2}_{i}\left(\mathbf{\Sigma_{\mathcal{F}}O^{M}D}\right)=\sigma^{2}_{i}\left(\mathbf{P}\right)$, in the case of uncorrelated noise, this is equivalent to minimizing $\sum^{\left|F\right|}_{i}\sigma^{2}_{i}$, which will reduce the amplification of noise in the signal recovery process,
@@
6) Maximization of the minimum singular value of $\mathbf{D}\mathbf{O}^{-\mathrm{M}}_{|\mathcal{F}|}$ (MaxSigMin): The strategy is designed to leverage the 2-norm to minimize NMSE, which is equivalent to minimizing the largest singular value of $\mathbf{P}$. In other words, we aim to maximize the smallest singular value $\sigma_{\mathrm{min}}$ of $\mathbf{D}\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$ for every qualified $\mathbf{D}$, thereby reducing the impact of noise in the recovered signal,
@@
7) Maximization of the singular value of $\mathbf{\mathbf{D}\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}}$ (MaxSig): The strategy aims to select columns of the matrix $\mathbf{O}^{-\mathrm{M}}_{|\mathcal{F}|}$ to maximize its singular values. Although this strategy is not directly related to NMSE optimization, it is easy to implement and demonstrates good performance in practice, improving the condition number of the sampling matrix and thus enhancing the energy concentration of the signal,
@@
The maximization of signal energy concentration in 2) and 3) is equivalent to ensuring perfect localization of the graph signal in the vertex and spectral domains, as described by Eqs. 12 and 13.
Remark 5.
Given $\mathbf{x=B^{M}y}$, where $\mathbf{y}$ is fixed, maximizing signal energy concentration is tantamount to maximizing $|\mathbf{DB^{M}}|^{2}$ or $|\mathbf{O^{M}DB^{M}}|^{2}$. Since the operator matrix $\mathbf{B^{M}DB^{M}}$ is a positive semi-definite, 2) and 3) are approximately equivalent to the A-optimal design and D-optimal design [44].
By achieving perfect localization of signals in both vertex and spectral domains, it implies that the signals primarily concentrate on specific parts or nodes of the graph. Maximizing their energy on the sampling subset $\mathcal{S}$ ensures that the sampled signals encompass the main local structures and features of the graph, thereby capturing the essential information and dynamics of the signals more effectively. Thus, the perfect localization of signals in vertex and spectral domains is closely intertwined with maximizing signal energy on the sampling subset $\mathcal{S}$. We summarize the A, D, E, T-optimal designs related to the seven sampling strategies along with their objective functions in Table 1.
@@
A-optimal $\min\mathrm{tr}\left[\left(\mathbf{B^{M}DB^{M}}\right)^{-1}\right]$ $\max\left|\mathbf{Dx}\right|^{2}$ MaxVertex
D-optimal $\min\mathrm{det}\left[\left(\mathbf{B^{M}DB^{M}}\right)^{-1}\right]$ $\max\left|\mathbf{O^{M}Dx}\right|^{2}$ MaxSpec
@@
As the graph signals are all $\mathbf{M}$-bandlimited after satisfying the localization condition, i.e., $\mathbf{x}\in\mathrm{BL}_{|\mathcal{F}|}(\mathbf{O^{M}})$, in other words, for the vertex set $\mathcal{V}$, its sampling subset $\mathcal{S}$ contains all the features and information of its graph signal. It satisfies the property of submodular functions, meaning that the marginal benefit of adding locally decreases with the increase of selected elements. More specifically, when we add an element to subset $\mathcal{S}$, although it may increase the number of covered elements, the marginal contribution of each newly added element gradually diminishes as the selected elements increase. Therefore, to address the sampling strategies mentioned above, we employ a greedy algorithm to tackle this selection challenge, aiming to determine the optimal sampling operator. Greedy algorithms have been shown to closely approximate the global optimum when solving similar matrix approximation optimization problems [45]. Since strategies 1), 2), and 7) are essentially equivalent, as are strategies 3) and 4), and strategies 5) and 6), we present Algorithm 1, which summarizes the three sampling strategies 4), 6), and 7).
  Input: $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$ (first $|\mathcal{F}|$ columns of $\mathbf{O}^{-\mathbf{M}}$), $|\mathcal{S}|$ (number of samples), $N$ (number of rows in $\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}$)
  Output: $\mathcal{S}$ (sampling set)
  for $k=1:|\mathcal{S}|$ do
     $s=\arg\max_{j\in\{1:N\}-\mathcal{S}}\sum^{|\mathcal{F}|}_{i=1}\sigma^{2}_{i}\left(\left(\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}\right)_{\mathcal{S}\cup\{j\}}\right)$
      or $s=\arg\max_{j\in\{1:N\}-\mathcal{S}}\sigma_{\mathrm{min}}\left(\left(\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}\right)_{\mathcal{S}+\{j\}}\right)$
      or $s=\arg\max_{j\in\{1:N\}-\mathcal{S}}\prod^{|\mathcal{F}|}_{i=1}\lambda_{i}\left(\left(\left(\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}\right)^{\ast}_{\mathcal{S}\cup\{j\}}\left(\mathbf{O}^{-\mathbf{M}}_{|\mathcal{F}|}\right)_{\mathcal{S}\cup\{j\}}\right)\right)$
     $\mathcal{S}=\mathcal{S}+\{s\}$
  end for
  return  $\mathcal{S}$
We compared our proposed seven sampling strategies with random sampling under the GLCT sampling framework. As shown in Fig. 5, the NMSE is reported, defined as the error between the original and recovered signals divided by the original signal for each node. (a) is based on a ring graph with N = 32 nodes, and the signal is a $\mathbf{M}$-bandlimited signal with parameters $\alpha=0.8,\beta=32,\xi=0.5k+1,k=1,...,N$. The frequency bandwidth is $|\mathcal{F}|=4$. (b) is based on a swiss roll graph with N = 256 nodes, and the signal is a $\mathbf{M}$-bandlimited signal with same parameters. The frequency bandwidth is $|\mathcal{F}|=32$. We compare the NMSE of eight sampling methods as the number of sampled nodes increases for two different graph signals. We observed that, with a small number of sampling nodes, all methods except random sampling could effectively recover the signal.